# Toward building next-generation Geocoding systems: a systematic review

### v1

@@ 4.1 Geocoding Workflow
Figure 1 illustrates the geocoding workflow, which can be broadly divided into two main parts: geocoding textual retrieval and geocoding interpolation. In the textual retrieval phase, a geocoding system segments the input description to create a query string, searches reference datasets, and ranks the matched candidates based on their matching scores. Therefore, we further divide this phase into three components: (1) data storage, (2) input parsing, and (3) retrieval. These three components in geocoding textual retrieval phase are closely linked: the retrieval algorithm dictates how reference data should be segmented and stored, and how input address description should be parsed. In contrast, geocoding interpolation is a standalone component, focusing solely on deriving the final output coordinates using the geometry of retrieved reference data. In the following sections, we elaborate on the current research status of each component that formalize a geocoding system.
@@ 4.2 Geocoding Textual Retrieval
Table 3 summarizes the variations of three components (i.e., data storage, input parsing, and retrieval) in the geocoding textual retrieval phase as observed in the existing research literature.
@@ 4.2.1 Data storage
Data storage in existing geocoding systems can be summarized from three perspectives: segmentation type, data representation type, and storage media type.
Segmentation type refers to how we divide an entire address description record into several blocks for further processing and storage in the database. Here, we only provide a high-level abstraction of how segmentation of an address record is determined, as some string similarity algorithms—discussed later in this section—may further refine certain portions of the address segmentation. At a high level, segmentation types can be categorized into address component-based and non-address component-based types. Address component-based segmentation divides each address record according to standard address components, as illustrated in Figure 2. In contrast, non-address component-based segmentation allows for the combination of these standard components. The choice of segmentation type depends on the retrieval methods employed in a geocoding system, which will be discussed in detail in Section 4.2.3. For example, if a geocoding system utilizes the Per-attribute Score Ranking method [32], which gives each address component a weighting score, each address record is more likely to be segmented according to individual address components. If combining house number, street base name, and street type can improve the quality of retrieved matching candidates, these three address components should be merged, necessitating the use of non-address component segmentation. In addition to segmenting address records, studies have introduced new fields to improve retrieval quality. For example, alias fields for city and state have been incorporated to account for abbreviations or outdated names commonly found in short messages  [118].
@@
In the phonetic-based group, address descriptions are converted into alphanumeric-encoded strings using phonetic algorithms to handle misspellings based on pronunciation [97]. A commonly used algorithm in geocoding systems is Soundex [113, 32]. However, it is limited to English and sensitive to the position of typos [113]. To address these issues, the Double-Metaphone algorithm was introduced to accommodate non-English languages [80, 93]. However, it still faces challenges, such as different words sharing the same encoding [66].
The string-based group represents each address record as a string and employs string-based similarity algorithms to identify matching candidates. Various string-based similarity algorithms have been used in geocoding systems [81, 107, 87]. Depending on the particular algorithm, strings in each field may undergo modifications. For example, before calculating similarity, token-based algorithms such as q-gram algorithms [60] first break strings into smaller segments, a pad version of q-gram algorithm [53] add special padding characters to original strings. These string-based algorithms are typically used to match street-level descriptions and city names. Recently, Yin [107] applied q-gram methods to postal codes, outperforming the traditional approach of matching only the first three digits. Thus, the ultimate representation of each address record is closely linked to the choice of string-based similarity algorithms for each address field.
The dense vector-based representation is an emerging string encoding approach, driven by recent advances in NLP and vector databases [73, 98, 41]. This method converts address descriptions into dense vectors, effectively capturing semantic relationships between words to mitigate potential semantic errors in input data. Existing geocoding literature has explored their retrieval performance of various word embedding techniques, ranging from Word2Vec to transformers [65, 26, 19].
@@
Storage media type refers to where and how reference datasets are stored, which can significantly impact the implementation of a geocoding system. Different databases offer varying levels of support for data models and query languages, influencing system performance and functionality. However, limited literature have conducted experiments on the comparison of geocoding performance resulting from a database or particular implementation. One reason could be proprietary concerns, especially for commercial geocoding systems. In this section, we provide a brief overview of storage media type used by geocoding systems, encouraging further research to explore this area in more depth. Historically, SQL-based databases have been dominant in geocoding fields  [30, 104]. However, the key advantage of SQL-based databases—their ability to join fields from different tables based on specific conditions—is primarily used for processing reference datasets rather than querying matching candidates  [29]. To this end, recent researches has examined the scalability of NoSQL based database, with the document-oriented and easy data partitioning design  [17, 54, 40]. Since geocoding involves searching within a reference dataset, which is treated as a fixed resource pool, some geocoding systems convert the entire reference dataset into a binary file for more efficient search processing  [^2] https://help.precisely.com/r/Spectrum/global_geocoding/23.1/en-US/Spectrum-Global-Geocoding-Guide/Custom-Dataset-Builder/Building-a-Custom-Dataset-for-Geocoding-and-Typeahead. Due to potential commercial interests, limited research has been published on the performance of geocoding systems that utilize binary files as their reference dataset format. Establishing benchmarks for geocoding performance across different storage media could provide valuable insight for developing next-generation systems. Furthermore, the growing adoption of dense vectors for string matching and retrieval highlights a promising research direction—exploring the application of vector databases in geocoding.
4.2.2 Input address parsing
@@
Geocoding address parsing (referred to as geocoding parsing) is a key focus in geocoding research due to its critical role in shaping how the query string is constructed and, ultimately, influencing the quality of output geocodes, in line with the geocoding “garbage in, garbage out” principle. This process involves the segmentation of input location descriptions and labeling of each token  [4]. The specific tokens used should depend on the geocoding system’s retrieval algorithm, namely how the reference data is segmented and how match candidates are ranked. Standard address components, as shown in Figure 2, are commonly used as target labels assigned to input address descriptions. It’s important to distinguish geocoding parsing from more general address parsers, which focus on identifying and extracting location descriptions from text and assigning entity labels (e.g., LOC). This broader task, known as toponym recognition or text geocoding, doesn’t require parsing the finer elements of addresses. Recent surveys by Hu et al [45] and Melo and Martins [72] provide comprehensive reviews on toponym recognition. However, to the best of the authors’ knowledge, few surveys have been published that thoroughly and specifically review geocoding address parsing techniques.
@@
Existing geocoding parsing solutions can be categorized into three approaches: (1) rule-based, (2) statistical-based, and (3) neural network-based. Compared to statistical-based and neural network-based approaches, rule-based approaches take a “manual” approach, as they utilize the structure and hierarchy of defined address system schemas to create logic to assign labels to input address descriptions  [31]. For example, in the U.S. address system, transitions between address components- house numbers followed by street level descriptions and 5-digit postal code typicall appears at the end of address description, are predictable. Current research on rule-based address parsers focuses on (1) abstracting address system hierarchies and (2) applying search algorithms (e.g., sliding window and heuristic search) and string-matching methods (e.g., edit distance) along with gazetteers to process addresses from various countries, including China, U.S., Turkey, and Japan  [119, 29, 67, 69]. However, limitations of rule-based methods have also been pinpointed. Address systems differ across countries, cultures, and languages, requiring a unique abstraction of each country’s hierarchy. In countries without a well-defined address system, rule-based methods are not applicable. Additionally, these methods rely on lexicons to recognize address components, making them prone to “Out of Vocabulary” issues when user inputs vary in quality or description  [62].
@@
To address these challenges, both statistical and neural network-based approaches have been developed, utilizing annotated training data to “learn” transitions between possible address components. Among the statistical methods, two popular models—Hidden Markov Models (HMMs) and Conditional Random Fields (CRFs)—are frequently applied for address parsing tasks [14, 94]. HMMs aim to maximize the joint probability between the observed sequence and states, assuming independence between each word and state, while CRFs improve on this by maximizing the conditional probability of labels given an observation sequence, using undirected graphical models. Initially used for NER tasks, both models have been extended to annotate address components. While most studies evaluate their performance across various countries’ address formats, further research seeks to enhance performance by integrating address system rules [62], stochastic regular grammar-based scoring functions [101], and by reducing the amount of training data required [20].
@@
Since address descriptions can be viewed as a sequence of data (i.e., text), it falls into the scope of Recurrent Neural Networks (RNNs), which aim to consider both the input for the current elements and preceding elements for each element in sequences. Among these RNN model architectures, Long Short-term Memory Networks (LSTMs), which introduces a memory-cell to capture long-range dependencies  [44], has shown advantages over traditional RNN models. Variations of LSTM models, including Bidirectional LSTM, attention-based LSTM, and hybrid LSTM-CRF models, have been developed and evaluated in NER tasks [57, 121, 12]. These LSTM-based architectures have inspired significant research into place name (LOC tag) recognition in various contexts [100, 6]. As noted earlier in this section, while address labeling and place name recognition both fall within the NER field and share certain similarities, their contextual sequences differ. Consequently, the performance of these models may not directly transfer to address labeling recognition and might require additional fine-tuning or evaluation. Although few studies have explicitly evaluated sequential neural networks for address labeling, these models have shown strong results. For example, [105] used subword embeddings and an LSTM-based embedding layer with a sequence-to-sequence (Seq2Seq) model to build an address parser, achieving state-of-the-art performance across addresses from 20 countries.
@@
More recently, LLMs, such as BERT [23] and GPT (Generative Pre-trained Transformer) [5], have revolutionized the field of NLP. These LLMs are pre-trained on vast amounts of unstructured text data, enabling them to learn rich linguistic representations and capture complex language patterns. One way to fine-tuning LLMs to the specific address parsing task. For example, [63] proposed a BERT-based model for address parsing called AddressBERT. They fine-tuned BERT on a large dataset of labeled address strings and achieved state-of-the-art performance in parsing addresses from multiple countries and languages. The model learns to identify and extract relevant address components based on the learned patterns and context. The other way is to leverage prompt engineering with generative AI. For instance, [111] benchmarked the performance of ChatGPT, with a specific prompt with just three address labels, using 230,000 U.S. address descriptions containing various input errors. The study showed that ChatGPT produced promising parsing results, compared to LSTM-based address parsers. Similarly, [48] employed a prompt-based approach using ChatLM to parse Chinese addresses. However, most of existing work focus on processing postal address alone, it would be more interesting to develop and evaluate an LLM-based parser that is capable of recognizing location descriptions embedded within more complex contexts, such as colloquial short-text sentences.
@@ 4.2.3 Retrieval
[…]
@@
[32] Phonetic and string similarity Predefined weight Deterministic
[64] Phonetic and string similarity Probabilistic weight Probabilistic
[19] Vector similarity Random Forest + XGBoost Machine learning
[82] String similarity SVM Machine learning
[65] Vector similarity ESIM Machine learning
[25] Pre-train model embedding K-NN Similarity Machine learning
It’s worthy noting that all these similarity methods and classifier could be combined randomly. The detail of these similarity methods have been discussed in Section 4.2.1, as the choice of similarity calculation methods also has impacts on how data is segmented and stored in a database. Thus, in this section, we focus on classifiers utilized by different geocoding systems, as existing literature used it to determine the type for a geocoding system. Existing literature [4, 31] mainly split the classifier into two groups: deterministic and probabilistic. We define a new category of machine learning for geocoding systems that employ a machine learning based classifier.
A deterministic classifier typically uses a predefined scoring mechanism to determine if matched candidates are found and rank these candidates. For example, [32] proposed a Per-attribute Score Ranking method to evaluate geocode candidates by assigning scores based on the match status of individual address components, each pre-assigned a weight reflecting its importance in determining a match. In addition to these address component scores, the method requires a minimum matching score to distinguish between match and no-match statuses. These empirically-determined score machisms limit its applicability to (1) address descriptions from other address systems and (2) matching field that differ from the scoring field.
Unlike a deterministic classifier, which uses empirically determined weights for each matching field, a probabilistic classifier assigns weights proportional to the frequency of each field in the reference datasets. One classic weight calculation method is known as the Fellegi and Sunter model [27], which is widely used in the domain of record linkage as follows [88, 38].
@@
where $W_{i}$ denotes the weight for $i$th field, $\gamma_{i}$ denotes the compared field $\gamma_{i}$ agrees, $m$ denotes the probability of the compared field agrees given that it is a true match $M$ and $\mu$ denotes the probability of the compared field agrees given that it is not a true match $U$. In geocoding practice  [88, 4], $P(\gamma_{i}\mid M)$ is assumed to be very close to 1, and $P(\gamma_{i}\mid U)$ is typically approximated by the fraction of the number of records with the field $\gamma_{i}$ and the number of records in collections. Later on, researchers started to adopt full-text document search engines (e.g., Elasticsearch  [^3] https://www.elastic.co/elasticsearch) to build geocoding systems  [15, 54]. Thus, two weight calculation methods: Term Frequency - Inverse Document Frequency (Equation 3) and Best Match 25 (Equation 4), which are commonly used in IR, have been used as the classifier to rank geocoding match candidates.
@@
While TF-IDF and BM25 differ in their weight calculation methods, both prioritize terms that appear less frequently in reference datasets. In TF-IDF, the inverse document frequency (IDF) is calculated as $IDF=\log\frac{N_{D}}{1+n_{t}}$, where $N_{D}$ is the total number of documents in the corpus $D$, and $n_{t}$ is the number of documents containing the term $t$. In contrast, BM25 incorporates the absence of the term in the corpus and applies smoothing as follows: $IDF=\log\frac{N_{D}-n_{t}+0.5}{n_{t}+0.5}$. Regarding term frequency, TF-IDF directly uses the raw term frequency without adjustment, while BM25 introduces saturation to limit the dominance of frequently occurring terms, thereby preventing them from disproportionately influencing the relevance score. While both methods have been utilized in production geocoding systems, no systematic performance comparison has been conducted using a unified benchmark dataset and standardized metrics.
A machine learning classifier learns thresholds to discriminate similarity values to each comparison on the field in an address description using annotated matched and unmatched address pairs using machine learning algorithms. As shown in Table 4, these algorithms range from classical machine learning classification model to neural network-based approaches. [19] combined Random Forest and XGBoost classifiers, using Word2Vec embeddings as input for matching U.K. addresses. This ensemble approach outperformed each classifier individually. [82] compared the performance of linear and nonlinear Support Vector Machines (SVMs) for address matching in Iran, demonstrating that nonlinear SVMs outperformed their linear counterparts. This result aligns with the findings from [19], which also reported superior performance of nonlinear classifiers compared to linear ones. With recent advances in NLP, [65] trained, using a corpus of Chinese addresses, a Word2Vec embedding model and an Enhanced Sequential Inference Model (ESIM) to predict match and unmatch labels for two addresses, outperforming traditional classifiers like Random Forest and SVM with the same Word2Vec embeddings. Instead of obtaining word embedding for segmented address components, [25] fine-tuned a pre-trained DistilBERT model to generate vector representations for a whole address description. They then retrieved the 10 most similar addresses and re-ranked them to derive the final output by processing these pairs through a transformer encoder. This approach achieved higher matching accuracy and shorter inference times compared to the BM25 method for matching Portuguese addresses.

### v2

@@ 4.1 Geocoding Workflow Decomposition
Figure 1 illustrates the geocoding workflow, which can be broadly divided into two main parts: geocoding textual retrieval and geocoding interpolation. The goal of the geocoding textual retrieval is to identify the record in the reference dataset that most closely corresponds to the input location description, whereas the geocoding interpolation derives the final output geocodes from the retrieved records based on their spatial relationships. During the textual retrieval phase, the geocoding system parses the input description into system-interpretable components, constructs a query string, searches the reference datasets, and ranks the retrieved candidates according to their matching scores. Therefore, we further divide this phase into three closely linked components: (1) data storage, (2) input parsing, and (3) retrieval, as the retrieval algorithm dictates how reference data should be segmented and stored, and how input address description should be parsed. The second part: geocoding interpolation can be considered a standalone component that focuses solely on deriving the final output coordinates using the geometry and spatial relationship of the retrieved reference data. Based on this decomposition, we now elaborate on the current state of each of these four components in the following sections.
@@ 4.2 Geocoding Textual Retrieval
As discussed earlier, data storage, input parsing, and retrieval are tightly coupled within the textual retrieval phase. Table 4 summarizes the variations in these three components, grouped by their implementation methodologies, reported in existing research. We examine each of these variations in greater detail below.
@@ 4.2.1 Data storage
Data storage, functioning much like a digital gazetteer that the system consults, determines how the reference dataset is organized, accessed, and retrieved. It can be characterized from three perspectives: segmentation type, data representation type, and storage media.
Segmentation type refers to how we divide an entire address description record into several blocks for further processing and storage in the database. Here, we only provide a high-level abstraction of how segmentation of an address record is determined, as some string similarity algorithms—discussed later in this section—may further refine certain portions of the address segmentation. At a high level, segmentation types can be categorized into address component-based and non-address component-based types. Address component-based segmentation divides each address record based on standard address components defined by a country’s address system. Figure 2 presents an example of the standard address components defined in the United States. In contrast, non–address-component-based segmentation permits arbitrary combinations of standard components. For instance, forming a street-level description from a pre-direction, base street name, road type, and post-direction. The choice of segmentation type is controlled by the retrieval methods employed in a geocoding system, which will be discussed in detail in Section 4.2.3. For example, if a geocoding system utilizes the Per-attribute Score Ranking method [goldberg2010candidateselectioncriteria], which gives each address component a weighting score, each address record is more likely to be segmented according to individual address components. If combining house number, street base name, and street type can improve the quality of retrieved matching candidates, these address components then should be merged, necessitating the use of non-address component segmentation. In addition to standard address components, studies have introduced new fields to improve retrieval quality. For example, alias fields for city and state have been incorporated to account for abbreviations or outdated names commonly found in short messages  [zhang2014geocodingTwitter].
@@
In the phonetic-based group, address descriptions are converted into alphanumeric-encoded strings using phonetic algorithms to handle misspellings based on pronunciation [toutanova2002pronunciation]. A commonly used algorithm in geocoding systems is Soundex [zandbergen2008comparison, goldberg2010candidateselectioncriteria]. However, it is limited to English and sensitive to the position of typos [zandbergen2008comparison]. To address these issues, the Double-Metaphone algorithm was introduced to accommodate non-English languages [philips2000doublemetaphone, shresthaa2011DMGeocoding]. However, it still faces challenges, such as different words sharing the same encoding [lisbach2013linguistic].
The string-based group represents each address record as a string and employs string-based similarity algorithms to identify matching candidates. Various string-based similarity algorithms have been used in geocoding systems [Ranzijn2013AddressMatchingTechniques, yin2021systematic, santos2018learning]. Depending on the particular algorithm, strings in each field may undergo modifications. For example, before calculating similarity, token-based algorithms such as q-gram algorithms [li200qGramFilterByThreshold] first break strings into smaller segments, a pad version of q-gram algorithm [keskustalo2003PadVersionQGramIsBetter] add special padding characters to original strings. These string-based algorithms are typically used to match street-level descriptions and city names. Recently, yin2021systematic applied q-gram methods to postal codes, outperforming the traditional approach of matching only the first three digits. Thus, the ultimate representation of each address record is closely linked to the choice of string-based similarity algorithms for each address field.
The dense vector-based representation is an emerging string encoding approach, driven by recent advances in NLP and vector databases [mikolov2013efficient, vaswani2017attention, han2023comprehensive]. This method converts address descriptions into dense vectors, effectively capturing semantic relationships between words to mitigate potential semantic errors in input data. Existing geocoding literature has explored their retrieval performance of various word embedding techniques, ranging from Word2Vec to transformers [lin2020W2VforGeocoding, duarte2023transformergeocoding, comber2019geocodingMatchWord2VecCDF].
@@
Storage media type refers to where and how reference datasets are stored, which can significantly impact the implementation of a geocoding system. Different databases offer varying levels of support for data models and query languages, influencing system performance and functionality. However, limited literature have conducted experiments on the comparison of geocoding performance resulting from a database or particular implementation. One reason could be proprietary concerns, especially for commercial geocoding systems. In this section, we provide a brief overview of storage media type used by geocoding systems, encouraging further research to explore this area in more depth. Historically, SQL-based databases have been dominant in geocoding fields  [Goldberg2011ImprovingMatchRateVaryingBlock, Xu2012GeocodingBillionAddress]. However, the key advantage of SQL-based databases—their ability to join fields from different tables based on specific conditions—is mainly used to process reference datasets rather than to query matching candidates  [goldberg2008geocodingbest]. To this end, recent researches have examined the scalability for NoSQL-based database, with the document-oriented and easy data partitioning design  [clemens2020geocodingPhdDissertation, knoblock2017ElasticSpatialExampleGeocoding, guo2020geocodingDatabase]. Since geocoding involves searching within a reference dataset, which is treated as a fixed resource pool, some geocoding systems convert the entire reference dataset into a binary file for more efficient search processing  [^2] https://help.precisely.com/r/Spectrum/global_geocoding/23.1/en-US/Spectrum-Global-Geocoding-Guide/Custom-Dataset-Builder/Building-a-Custom-Dataset-for-Geocoding-and-Typeahead. Due to potential commercial interests, limited research has been published on the performance of geocoding systems that utilize binary files as their reference dataset format. Thus, establishing benchmarks for geocoding performance across different storage media could provide valuable insight about which storage media are the most efficient.
4.2.2 Input parsing
@@
Geocoding address parsing (referred to as geocoding parsing) is a key focus in geocoding research, as it governs the construction of the query string in the downstream geocoding workflow and therefore impacts the quality of the resulting geocodes. This process involves segmenting the input location description and assigning a label to each resulting token [boscoe2008science]. The specific tokens used should depend on the geocoding system’s retrieval algorithm, namely how the reference data is segmented and how match candidates are ranked. Standard address components, as shown in Figure 2, are commonly used as target labels assigned to input address descriptions. It’s important to distinguish geocoding parsing from more general address parsers, which focus on identifying and extracting location descriptions from text and assigning entity labels (e.g., LOC). This broader task, known as toponym recognition or text geocoding, doesn’t require parsing the finer elements of addresses. Recent surveys by hu2023location and melo2017SurveyOfGeocodingTextualDocuments provide comprehensive reviews on toponym recognition. However, to the best of the authors’ knowledge, few surveys have been published that thoroughly and specifically review geocoding address parsing techniques.
@@
Existing geocoding parsing solutions can be categorized into three approaches: (1) rule-based, (2) statistical-based, and (3) neural network-based. Compared to statistical-based and neural network-based approaches, rule-based approaches take a “manual” approach, as they utilize the structure and hierarchy of defined address system schemas to create logic to assign labels to input address descriptions  [goldberg2013geocodingTechniques]. For example, in the U.S. address system, transitions between address components- house numbers followed by street level descriptions and 5-digit postal code typicall appears at the end of address description, are predictable. Current research on rule-based address parsers focuses on (1) abstracting address system hierarchies and (2) applying search algorithms (e.g., sliding window and heuristic search) and string-matching methods (e.g., edit distance) along with gazetteers to process addresses from various countries, including China, U.S., Turkey, and Japan  [zhnag2010ruleBaseChineseAddress, goldberg2008geocodingbest, liu2002JapanLexicon, matci2018address]. However, limitations of rule-based methods have also been pinpointed. Address systems differ across countries, cultures, and languages, requiring a unique abstraction of each country’s hierarchy. In countries without a well-defined address system, rule-based methods are not applicable. Additionally, these methods rely on lexicons to recognize address components, making them prone to “Out of Vocabulary” issues when user inputs vary in quality or description  [li2018hybridChineseAddressSegmentions].
@@
To address these challenges, both statistical and neural network-based approaches have been developed, utilizing annotated training data to “learn” transitions between possible address components. Among the statistical methods, two popular models—Hidden Markov Models (HMMs) and Conditional Random Fields (CRFs)—are frequently applied for address parsing tasks [Christen2005hmmAddressParserInputErrors, sun2017chineseCDFNER]. HMMs aim to maximize the joint probability between the observed sequence and states, assuming independence between each word and state, while CRFs improve on this by maximizing the conditional probability of labels given an observation sequence, using undirected graphical models. Initially used for NER tasks, both models have been extended to annotate address components. While most studies evaluate their performance across various countries’ address formats, further research seeks to enhance performance by integrating address system rules [li2018hybridChineseAddressSegmentions], stochastic regular grammar-based scoring functions [wang2016CRFAddressParser], and by reducing the amount of training data required [craig2019BingmapParserConcatenateW2Vfeatures].
@@
Since address descriptions can be viewed as a sequence of data (i.e., text), it falls into the scope of Recurrent Neural Networks (RNNs), which aim to consider both the input for the current elements and preceding elements for each element in sequences. Among these RNN model architectures, Long Short-term Memory Networks (LSTMs), which introduce a memory-cell to capture long-range dependencies  [hochreiter1997LSTM], have shown advantages over traditional RNN models. Variations of LSTM models, including Bidirectional LSTM, attention-based LSTM, and hybrid LSTM-CRF models, have been developed and evaluated in NER tasks [lample2016LSTMCRf, zhou2016attentionLSTM, chiu2016bidirectionalLSTM]. These LSTM-based architectures have inspired significant research into place name (LOC tag) recognition in various contexts [wang2020neurotpr, cardoso2019using]. As noted earlier in this section, while address labeling and place name recognition both fall within the NER field and share certain similarities, their contextual sequences differ. Consequently, the performance of these models may not directly transfer to address labeling recognition and might require additional fine-tuning or evaluation. Although few studies have explicitly evaluated sequential neural networks for address labeling, these models have shown strong results. For example, [yassine2021universalAddressParsing] used subword embeddings and an LSTM-based embedding layer with a sequence-to-sequence (Seq2Seq) model to build an address parser, achieving state-of-the-art performance across addresses from 20 countries.
@@
More recently, LLMs, such as BERT [devlin2018bert] and GPT (Generative Pre-trained Transformer) [brown2020language], have revolutionized the field of NLP. These LLMs are pre-trained on vast amounts of unstructured text data, enabling them to learn rich linguistic representations and capture complex language patterns. One way to fine-tune LLMs to the specific address parsing task. For example, [li2023enex] proposed a BERT-based model for address parsing called AddressBERT. They fine-tuned BERT on a large dataset of labeled address strings and achieved state-of-the-art performance in parsing addresses from multiple countries and languages. The model learns to identify and extract relevant address components based on the learned patterns and context. The other way is to leverage prompt engineering with generative AI. For instance, [yin2023chatgpt] benchmarked the performance of ChatGPT, with a specific prompt with just three address labels, using 230,000 U.S. address descriptions containing various input errors. The study showed that ChatGPT produced promising parsing results, compared to LSTM-based address parsers. Similarly, [huang2024geoagent] employed a prompt-based approach using ChatLM to parse Chinese addresses.
@@ 4.2.3 Retrieval
$\textit{Matching score}=\textit{Classifier(Similarity)}$ (1)
@@
[goldberg2010candidateselectioncriteria] Phonetic and string similarity Predefined weight Deterministic
[li2014region] Phonetic and string similarity Probabilistic weight Probabilistic
[comber2019geocodingMatchWord2VecCDF] Vector similarity Random Forest + XGBoost Machine learning
[rezaei2024address] String similarity SVM Machine learning
[lin2020W2VforGeocoding] Vector similarity ESIM Machine learning
[duarte2023improving] Pre-train model embedding K-NN Similarity Machine learning
It’s worthy noting that all these similarity methods and classifier could be combined randomly. The details of these similarity methods have been discussed in Section 4.2.1, as the choice of similarity calculation methods also has impacts on how data are segmented and stored in a database. Thus, in this section, we focus on classifiers utilized by different geocoding systems, since the existing literature often used the type of classifier to differentiate the types of geocoding systems: deterministic and probabilistic [boscoe2008science, goldberg2013geocodingTechniques]. We define a new category of machine learning for geocoding systems that employ a machine learning based classifier.
A deterministic classifier typically uses a predefined scoring mechanism to determine if matched candidates are found and then rank these candidates. For example, [goldberg2010candidateselectioncriteria] proposed a Per-attribute Score Ranking method to evaluate geocode candidates by assigning scores based on the match status of individual address components, each pre-assigned a weight reflecting its importance in determining a match. In addition to these address component scores, the method requires a minimum matching score to distinguish between match and non-match statuses. These empirically-determined score mechanisms limit its applicability to (1) address descriptions from other address systems and (2) matching fields that differ from the scoring field.
Unlike a deterministic classifier, which uses empirically determined weights for each matching field, a probabilistic classifier assigns weights proportional to the frequency of each field in the reference datasets. A classic weight calculation method is known as the Fellegi and Sunter model [fellegi1969theoryFSmodel], which is widely used in the domain of record linkage as follows [sayers2015probabilisticRecordLinkageReview, gu2003recordLinkageReview].
@@
where $W_{i}$ denotes the weight for the $i$th field, $\gamma_{i}$ denotes the agreement of the compared field $\gamma_{i}$, $m$ denotes the probability that the compared field agrees given that it is a true match $M$ and $\mu$ denotes the probability that the compared field agrees given that it is not a true match $U$. In geocoding practice  [sayers2015probabilisticRecordLinkageReview, boscoe2008science], $P(\gamma_{i}\mid M)$ is assumed to be very close to 1, and $P(\gamma_{i}\mid U)$ is typically approximated by the fraction of the number of records with the field $\gamma_{i}$ and the number of records in collections. Later, researchers started to adopt full-text document search engines (e.g., Elasticsearch  [^3] https://www.elastic.co/elasticsearch) to build geocoding systems  [clemens2015geocodingESwithAddrTokenDropped, knoblock2017ElasticSpatialExampleGeocoding]. Thus, two weight calculation methods: Term Frequency - Inverse Document Frequency (Equation 3) and Best Match 25 (Equation 4), which are commonly used in IR, have been used as the classifier to rank geocoding match candidates.
@@
While TF-IDF and BM25 differ in their weight calculation methods, both prioritize terms that appear less frequently in reference datasets. In TF-IDF, the inverse document frequency (IDF) is calculated as $IDF=\log\frac{N_{D}}{1+n_{t}}$, where $N_{D}$ is the total number of documents in the corpus $D$, and $n_{t}$ is the number of documents that contain the term $t$. In contrast, BM25 incorporates the absence of the term in the corpus and applies smoothing as follows: $IDF=\log\frac{N_{D}-n_{t}+0.5}{n_{t}+0.5}$. Regarding term frequency, TF-IDF directly uses the raw term frequency without adjustment, while BM25 introduces saturation to limit the dominance of frequently occurring terms, thereby preventing them from disproportionately influencing the relevance score.
A machine learning classifier learns thresholds to discriminate similarity values to each comparison on the field in an address description using annotated matched and unmatched address pairs using machine learning algorithms. As shown in Table 5, these algorithms range from classical machine learning classification models to neural network-based approaches. [comber2019geocodingMatchWord2VecCDF] combined Random Forest and XGBoost classifiers, using Word2Vec embeddings as input to match U.K. addresses. This ensemble approach outperformed each classifier individually. [rezaei2024address] compared the performance of linear and nonlinear Support Vector Machines (SVMs) for address matching in Iran, demonstrating that nonlinear SVMs outperformed their linear counterparts. This result is in line with the findings of [comber2019geocodingMatchWord2VecCDF], which also reported superior performance of nonlinear classifiers compared to linear ones. With recent advances in NLP, [lin2020W2VforGeocoding] trained, using a corpus of Chinese addresses, a Word2Vec embedding model, and an Enhanced Sequential Inference Model (ESIM) to predict match and unmatch labels for two addresses, outperforming traditional classifiers like Random Forest and SVM with the same Word2Vec embeddings. Instead of obtaining word embedding for segmented address components, [duarte2023improving] fine-tuned a pre-trained DistilBERT model to generate vector representations for a whole address description. They then retrieved the 10 most similar addresses and re-ranked them to derive the final output by processing these pairs through a transformer encoder. This approach achieved higher matching accuracy and shorter inference times compared to the BM25 method to match Portuguese addresses.
To date, all three types of classifiers have been utilized in geocoding systems. To make an informed choice, different classifiers should be systematically evaluated on a unified data set that closely reflects the quality and characteristics of the inputs the system is expected to encounter in production.